OpenStudy (anonymous):

3(3x - 2) - 29 = 3x + 1

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

9x - 6 - 29 = 3x + 1 9x - 35 = 3x + 1 6x = 36 x = 6

OpenStudy (anonymous):

graph y= 39x - 35 and g = 3x + 1, their intersections will give you the solution